For i = the square root of -1, if 3i (2 + 5i) = x + 6i, then x = ?
Answer by fractalier(6550)

For 3i(2 + 5i) = x + 6i, we have
3i(2 + 5i) = 6i + 15i^2 = -15 + 6i
so x = -15